A buffer solution is prepared by mixing CH3COONa and CH 3COOH. The pk, of acetic acid is 4.74. To maintain the PH of the buffer solution as 6.04, the concentration ratio of CH3COONa and CH3COOH to be maintained is

Grade:12th pass

1 Answers

Devang
5 years ago
for buffer solution 
             pH = pKa + log [ salt ] / [Acid]        …..(1)
    where Ka = Dissociation constant of acid 
in this question pH = 6.04 ,  pKa  =  4.74 ,  [ salt ] / [Acid] = ?
 from eq. (1) 
            6.04 = 4.74 + log [ salt ] / [Acid]  
           1.3    =   log [ salt ] / [Acid]
 [ salt ] / [Acid]  = Antilog (1.3)
 [ salt ] / [Acid]  = 1.995 * 10 = 19.95